OpenStudy (lifeisadangerousgame):

Write the equation of the line in point-slope form. Show all work. (5 points) (3, 5); slope = -9 This is my answer. I'm horrible at slope and if it's wrong I'm not intentionally getting it wrong. I'm just really bad at slope :/ y - 5 = -9(x - 3) y = 27 + 9x + 5 y = 9x + 32

OpenStudy (anonymous):

y-5=-9(x-3) <------------That IS point slope form Point slope form of a line: \[y-y1=m(x-x1)\]

OpenStudy (anonymous):

In slope intercept form, \[y=mx+b\] y-5=-9x+27 y=-9x+32 You have a +9x in yours, not a -

OpenStudy (lifeisadangerousgame):

Ohh ok, thank you!

OpenStudy (lifeisadangerousgame):

Is it supposed to be y = 27 - 9x + 5 or do I have it right?

OpenStudy (anonymous):

in slope intercept form, y=-9x+32

OpenStudy (lifeisadangerousgame):

So wait it's supposed to be y - 5 = -9(x - 3) y = 27 - 9x + 5 y = 9x + 32 or y - 5 = -9(x - 3) y = - 9x + 32 y = 9x + 32

OpenStudy (anonymous):

Why are you removing the - sign from in front of the 9?

OpenStudy (lifeisadangerousgame):

It keeps removing it when I press post >:( It's there while I write it, then leaves when I press post. I meant y - 5 = -9(x - 3) y = -9x + 32 y = -9x + 32

OpenStudy (anonymous):

yeah that's right

OpenStudy (lifeisadangerousgame):

Thank you!!